OpenStudy (anonymous):

To solve the equation r plus six equals six, you should: subtract 0 from both sides. add 6 to both sides. subtract 6 from both sides. add twelve to both sides.

OpenStudy (kaylaprincess):

r + 6 = 6 r = 6 - 6 r = 0 Subtract 6

OpenStudy (anonymous):

@kaylaprincess Thank you!

OpenStudy (anonymous):

q